For series lcr circuit R=x L = 2 X c 
Impedance is??

Grade:12

2 Answers

Arun
4 years ago
Dear student
 
Impedence is = sqrt (XL^2 + Xc^2)
 
= sqrt (R^2 + R^2/4) = sqrt(5R^2/2) = R sqrt(5/2)
 
Hope it helps
 
Regards
Arun
Vikas TU
4 years ago
Dear student 
Impedance is Z = sqrt( R^2 + (Xl – Xc )^2 )
So, act accordingly . 
Z = sqrt(R^2 + (Xc )^2)
Hope this helps